OpenStudy (anonymous):

If a ¼ of a teaspoon is 1 ml, then how many milliliters are in 6 teaspoons? i know the answer is 24ml but for these types of questions that will pop up on my test.. what do i do if i don't know the conversions. for example this one is 1teas= 4ml.. i didnt know that till i looked it up but when i take the test i wont have that resource so is there another way to work this out not knowing 1teas=24ml

OpenStudy (anonymous):

i would say just try to memorize some conversions, though im assuming you dont have a photographic memory huh?

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

1/4 teaspoon = 1 ml 4*1/4 teaspoon = 4*1 ml ... Multiply both sides by 4. 1 teaspoon = 4 ml 6*1 teaspoon = 6*4 ml ... Multiply both sides by 6. 6 teaspoons = 24 ml

OpenStudy (anonymous):

my advice? obtain photographic memory XD lol nah i would try to remember the most used ones first, other than that i say write a conversion list on your palm XD

OpenStudy (anonymous):

in the second line why did u take times 4 to 1/4 and 1ml

jimthompson5910 (jim_thompson5910):

or you can do it this way 6 teaspoons = 6*(1 ml/(1/4 teaspoons)) 6 teaspoons = 6/(1/4) ml 6 teaspoons = 6*4 ml 6 teaspoons = 24 ml but this way may be a bit more complicated

OpenStudy (anonymous):

To get rid of the fraction.

jimthompson5910 (jim_thompson5910):

because I wanted the 1/4 teaspoons to turn into 1 teaspoon that way, I can easily multiply that 1 by 6 to get 6 teaspoons

OpenStudy (anonymous):

she already has the answer down, shes needing to memorize conversion formulas or needs an easy way to remember them correct?

OpenStudy (anonymous):

but when im sitting here doing the problem what tells me oh im going to take 4 times or 6 times why four and why six

jimthompson5910 (jim_thompson5910):

you start with 1/4 teaspoon = 1 ml you want something in the form 1 teaspoon = .... basically you want to know how many ml are in one teaspoon

OpenStudy (anonymous):

jim, so you used the 4 from 1/4 the denom?

jimthompson5910 (jim_thompson5910):

once you have 1 teaspoon = 4 ml you can multiply both sides by 6 to get 6 teaspoons = 24 ml

jimthompson5910 (jim_thompson5910):

yes multiplying 1/4 by 4 will give you 1

OpenStudy (anonymous):

ok so why my second step to take it all by six

jimthompson5910 (jim_thompson5910):

because we want to know how many ml are in 6 teaspoons

OpenStudy (anonymous):

haha ok duh

OpenStudy (anonymous):

so could you give me another example like if it wasnt 1/4 if it was like 3/4

OpenStudy (anonymous):

or 2/8

jimthompson5910 (jim_thompson5910):

ok if it was 3/4, then 3/4 teaspoons = 1 ml 3 teaspoons = 4 ml ... Multiply both sides by 4. 1 teaspoon = 4/3 ml ... Divide both sides by 3. Then you multiply both sides by 6 to get 6 teaspoons = 24/3 ml 6 teaspoons = 8 ml but this is of course 3/4 teaspoons = 1 ml is true

jimthompson5910 (jim_thompson5910):

if 3/4 teaspoons = 1 ml is true* i meant

jimthompson5910 (jim_thompson5910):

and 2/8 reduces to 1/4, so 2/8 = 1/4

OpenStudy (anonymous):

haha okay thank you.. ill have another question posted soon could u keep an eye out for me please ur helpful than most